OpenStudy (anonymous):

write a rule for a linear function in the table. x f (x) 1 -7 2 -10 3 -13 4 -16 A.) f (x)=-1/3x =-4 B.) f (x)=3x+4 C.) f (x)=-3x -4 D.) f (x)= x-7 can somebody help me with this?

OpenStudy (anonymous):

put the values of x in equations

OpenStudy (anonymous):

idk how to do that. im not good at math

OpenStudy (anonymous):

okay is A.) f (x)=-1/3x =-4 the option you gave correct? check and tell if u typed it wrong

OpenStudy (anonymous):

im stuck at the slope 3y/1x = im stuck right here